$SPAD/src/input e04jaf.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)clear all
showArrayValues true
showScalarValues true
n:=4
ibound:=0
liw:=6
lw:=54
bl:Matrix SF:=
 [[1 ,-2 ,-1.0e-6 ,1 ]]
bu:Matrix SF:=
 [[3 ,0 ,1.0e6 ,3 ]]
x:Matrix SF:=
 [[3 ,-1 ,0 ,1 ]]
ef:Expression Float:=
 ((XC[1]+10*XC[2])**2+5*(XC[3]-XC[4])**2+(XC[2]-2*XC[3])**4+_
   10*(XC[1]-XC[4])**4)::EXPR FLOAT
funct1:Asp24(FUNCT1):=retract ef
result:=e04jaf(n,ibound,liw,lw,bl,bu,x,-1,funct1)

———-

References^{[1] nothing}

4